# Arena: A Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

### v1

@@ 4. Implementation
We develop a prototype of Arena on an edge server equipped with two Intel(R) Xeon(R) Silver 4310 CPUs, 128 GB of RAM, and two NVIDIA GeForce RTX 4090 GPUs (16384 CUDA Cores) , each with 24 GB of VRAM. An NVIDIA Jetson AGX Orin running Ubuntu 22.04 LTS serves as the camera. The edge server and camera are connected with the TP-LINK TL-WDR5620 through a 2.4GHz WiFi network. The bandwidth is 93.9 Mbps, measured by iperf. Arena performs preprocessing on the camera to obtain the PoIs that need to be transmitted. The camera utilizes the HTTP protocol and creates request sessions to send POST requests to the server. The server returns the position information for PoIs of the next frame to the camera in JSON format. The ViT detector is implemented with Python and trained using MMTracking (Contributors, 2020) framework on the edge server. The PPS is implemented with C++ on the camera. Figure 5 shows our experimental platform.
@@
Figure 5. Snapshot of our experimental evaluation hardware platform, including an edge server with GPUs and an NVIDIA Jetson AGX as the camera.
5. Evaluation
@@
5.1. Experimental Setup
5.1.1. Performance Metrics
@@ 5.1.2. Datasets
$\bullet$ MOT17-Det (Milan et al., 2016): It contains moving or stationary street views with pedestrian bounding box labels from indoor and outdoor settings across 11 distinct regions. The resolutions of these videos are either $640\times 480$ or $1920\times 1080$. We split the initial 75% of each video sequence to constitute our training set, with the remaining 25% serving as the validation set.
$\bullet$ AI City Challenge 2022 (AIC22) (Naphade et al., 2022): AIC22 is a real-world multi-camera traffic dataset including 42,683 frames and 88,820 vehicle bounding box labels. It contains video footage from a total of 59 surveillance cameras situated at five traffic intersections. We select videos from three intersections to constitute our training set while setting the videos from the remaining intersections as the validation set.
5.1.3. Arena’s Configurations
We employ Faster R-CNN (Ren et al., 2015) as the two-stage detector framework and ViT-small as the backbone, which features a 12-layer transformer encoder and $D=384$. MFR is a one-layer transformer decoder that comprises merely 2.07M (5% of total) parameters and requires approximately 17.16 (1.9% of total) GFlops. The backbone pre-trained weights are from (Ren et al., 2023). Unless specifically stated, the keyframe interval $K=5$, $p=0.9$, and $F=200$. For the expanded bbox margin $m$, we set it to $1$ for the MOT17Det dataset and $3$ for the AIC22 dataset.
5.1.4. Baselines
@@
$\bullet$ Faster R-CNN Detector (Ren et al., 2015) (FRD): Similar to FD, but the edge server employs a Faster R-CNN detector with ResNet-50.
@@
$\bullet$ DDS (Du et al., 2020): DDS transmits low-quality videos to the edge server to identify RoIs, which are fed back to the camera. The camera subsequently transmits these regions in high quality for second-round inference. We simulate different qualities by adjusting the resolution.
@@
$\bullet$ Reducto (Li et al., 2020): Reducto filters out frames with minimal impact on the accuracy by comparing low-level video feature differences between two adjacent frames at the camera to reduce bandwidth consumption. We employ Edge feature and utilize training sets to profile the optimal threshold.
@@
$\bullet$ ELF (Zhang et al., 2021a): ELF uses attention-based LSTM to predict the bounding box shifts frame by frame. It then employs a frame partitioning method to offload the region proposals to multiple edge servers. We utilize the official implementation and configure it for three servers with parallel offloading and inference.
5.2. Overall Performance of Arena
Figure 6. The normalized bandwidth usage of different methods on two datasets.
@@ 5.2. Overall Performance of Arena
Figure 7. The accuracy of different methods on two datasets. Arena can maintain accuracy losses within 1% and 4%.
@@
Figure 8. The average end-to-end latency per frame of different methods on two datasets. End-to-end latency includes a breakdown of preprocessing, transmission, and inference time.
@@
In Figure 6 and Figure 7, we report the normalized bandwidth usage and accuracy of different methods on two datasets. Figure 8 displays the end-to-end latency of each method, including a breakdown of preprocessing, transmission, and inference time. Among these baselines, Arena achieves the best trade-off among bandwidth, latency, and accuracy in both datasets. FD and FRD always transmit the full frame for inference, representing the upper bound for bandwidth consumption. They achieve the highest accuracy on their models but also result in the longest end-to-end latency. DDS achieves high bandwidth efficiency by transmitting only a few regions in high quality but inevitably incurs additional delay due to two rounds of inference. Reducto struggles to adapt to object detection tasks, as removing too many frames significantly reduces its accuracy. ELF maintains low bandwidth consumption and end-to-end latency, but this is contingent upon the premise of parallel offloading and inference by three servers. In contrast, Arena substantially saves bandwidth usage by pruning patches in non-keyframes, thus reducing both transmission delay and the model inference time. Meanwhile, Arena utilizes historical frame information to reconstruct features, minimizing the impact on accuracy. Specifically, Arena achieves a maximum acceleration of end-to-end latency by 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × in two datasets, using only 54% and 34% of the bandwidth, while maintaining accuracy losses within 1% and 4% as compared to FD, respectively.
@@
Figure 9. Visualization of Arena on two videos. In these two scenes, with a frame interval of 5, $m$ is set to 1 and 3 for MOT17 and AIC22, respectively, $p=0.9$, and $F=200$. Only the red patches in non-keyframe are used for transmission and inference.
@@
Figure 10. Heatmaps of patches identified as PoIs, where darker areas indicate a higher frequency of offloading to the edge server.
@@
Moreover, Figure 11 presents the cumulative distribution function (CDF) charts of the proportion of PoI extraction in non-keyframes across various scenes. Taking the AIC22 dataset as an example, PoIs are extracted in a more granular and sparse manner, with about 80% of non-keyframes requiring the transmission of less than 25% area of the frame. This is consistent with our findings in the motivation (Sec. 2.2). The results above reveal that Arena can effectively exploit the extensive redundancy present in the video to achieve end-to-end acceleration.
Figure 11. The cumulative distribution function (CDF) of PoI proportion of all non-keyframes.
5.3. Accuracy-Bandwidth Usage Trade-off
In the design of Arena, there is a trade-off between accuracy and bandwidth usage. As depicted in Figure 12, we observe that as the keyframe interval increases, there is a reduction in bandwidth usage, but there is also a decrease in accuracy. A similar observation is made in Figure 13, where reducing the expanded bbox margin significantly decreases bandwidth usage. However, there is also a slight accuracy drop.
This trade-off relationship provides us with insights for identifying the optimal configuration. We refer to these configurations, where accuracy and bandwidth usage are well balanced, as the Sweet Spot(s). For instance, in Figure 12, these sweet spots (e.g., $K=3,4,5$) exhibit a notable reduction in bandwidth usage without a significant decrease in accuracy, thus achieving an ideal balance. Specifically, when $K=3$, Arena achieves a 38.66% reduction in bandwidth with only 0.1% loss in accuracy for MOT17Det, and a 55.76% bandwidth saving at the expense of 2.5% accuracy loss for AIC22. Additionally, as shown in Figure 13, initially (e.g., $m=0,1,2$), both accuracy and bandwidth usage increase with the expanded bbox margin $m$. However, when further increases the expanded bbox margin $m$, the accuracy does not significantly improve, thereby reaching a sweet spot. Specifically, when $m=3$, the accuracy curve of Arena reaches a stable plateau, and at this point, the bandwidth usage is 60.56% and 33.67% of the original. These optimal configurations can be obtained through offline profiling or specific strategies, which we leave for future work.
@@ 5.3. Accuracy-Bandwidth Usage Trade-off
Next, we illustrate the impact of sampling rates within high probability regions on accuracy in Figure 14. For MOT17Det, accuracy remains relatively stable when the probability $p$ is within the range of 0.8 to 1.0. Similarly, for AIC22, a $p$ value exceeding 0.9 consistently keeps high accuracy. This indicates Arena’s strong capability to compensate for the input.
Figure 12. The impact of varying keyframe intervals on accuracy and bandwidth usage. ”Inf” denotes that all frames except the first one are considered as non-keyframes (i.e., $K=\infty$).
Figure 13. The impact of varying expanded bbox margin on accuracy and bandwidth usage.
Figure 14. The impact of different sampling rates on accuracy.
5.4. Model Analysis
@@
In this subsection, we compare the model complexity of using an ViT detector with other commonly used detectors. As shown in Table 2, we benchmark the GMACs and GFlops at a resolution of $800\times 1333$ using calflops [^1] https://github.com/MrYxJ/calculate-flops.pytorch and report the frame per second (FPS). ViT-small is employed as the backbone instead of the ViT-base version in Arena because it can achieve better real-time performance while already maintaining good accuracy. When processing the videos in the dataset, Arena operates with a computing load below 25% or 50% versions in the non-keyframe inference phase. This is because the proportion of PoI in most scenes is below 50% and even 25% as shown in Figure 11.
@@
Finally, we report the model precision in Table 3. Though Faster R-CNN offers the fastest inference speed, it requires the transmission of full frames, thus slowing down the end-to-end latency. Other models (Zhu et al., 2021; Zhang et al., 2022b) using the detection transformer (Carion et al., 2020) architecture, despite achieving commendable accuracy, fall short in real-time performance.
@@
Table 2. Comparison of model complexity with 800 $\times$ 1333 resolution. Here we compare the inference speed (FPS) and complexity (GFlops) between widely used object detectors with different ratios of patch proportion.
@@
Faster R-CNN (Ren et al., 2015) ResNet50 41.123 214.648 430.087 47.47
Deformable-DETR (Zhu et al., 2021) ResNet50 39.823 203.641 408.499 2.73
DINO (Zhang et al., 2022b) ResNet50 47.265 286.45 574.363 2.07
@@
Table 3. Model Precision
@@
Faster R-CNN (Ren et al., 2015) ResNet50 0.791 0.587 0.865 0.620
Deformable-DETR (Zhu et al., 2021) ResNet50 0.805 0.630 0.880 0.648
DINO (Zhang et al., 2022b) ResNet50 0.819 0.632 0.876 0.656
6. Related Work
6.1. Real-Time Video Analytics Systems
@@
Real-time video analytics is a computationally intensive task. Substantial effort focuses on enhancing the real-time performance of systems in various ways. For example, to reduce the transmission data, server-driven methods (Du et al., 2020; Zhang et al., 2022c, a) allow cameras to send low-quality videos to the edge server. The edge server then identifies RoIs and provides the camera with their position feedback. In the second transmission, only these RoIs, encoded in high quality, are sent to the server to be processed. Another line of work (Liu et al., 2019; Chen et al., 2022; Zhang et al., 2021a; Liu et al., 2022b) adopts a lightweight approach that directly extracts RoIs at the cameras and transmits only these regions to the edge server for inference. Additionally, some cloud-edge collaborative systems (Yang et al., 2023; Wang et al., 2022; Yi et al., 2020) enhance system efficiency by offloading partial computational tasks to the cloud. However, our work is the first to fully leverage characteristics of vision foundation models and achieve acceleration both on computing and transmission in the video analytics system.
6.2. Efficient Methods for Vision Transformer
@@
Utilizing the MSA mechanism in ViT leads to computational complexity that typically scales quadratically with the number of tokens. Consequently, recent efforts have focused on reducing the number of tokens to improve the efficiency of ViT. These methods primarily fall into two categories: token pruning (Rao et al., 2021; Liang et al., 2022; Long et al., 2023; Wei et al., 2023; Yin et al., 2022; Cao et al., 2023) and token merging (Marin et al., 2023; Zeng et al., 2022; Bolya et al., 2022).
@@
Just like what has been done to CNN-type architecture pruning in the past few years, token pruning is applied to accelerate ViT. For example, DynamicViT (Rao et al., 2021) introduces a trainable prediction module (i.e., MLP) that leverages both local and global token information to create a binary mask, which determines the token count for the next layer. Most merging-based methods rely heavily on clustering algorithms, such as K-Means (Marin et al., 2023), K-Medoids (Marin et al., 2023), and density-peak clustering with K-Nearest neighbours (Zeng et al., 2022) to cluster and merge tokens. Moreover, ToMe (Bolya et al., 2022) proposes a bipartite soft matching algorithm to cluster tokens and subsequently merge similar tokens, thus reducing the overall token count.
@@
These ViT acceleration techniques are not suited for Arena. First, they cannot reconstruct complete feature maps, which are necessary for dense prediction tasks. Second, they require the entire frame as input. Our proposed Arena, thinking out of the box, effectively adapts efficient ViT for video analytics with its advantages.
7. Conclusion

### v2

@@ IV Implementation
We develop a prototype of Arena on an edge server equipped with two Intel(R) Xeon(R) Silver 4310 CPUs, 128 GB of RAM, and two NVIDIA GeForce RTX 4090 GPUs (16384 CUDA Cores), each with 24 GB of VRAM. An NVIDIA Jetson AGX Orin running Ubuntu 22.04 LTS serves as the camera. The edge server and camera are connected with the TP-LINK TL-WDR5620 through a 2.4GHz WiFi network. The bandwidth is 93.9 Mbps, measured by iperf. Arena performs preprocessing on the camera to obtain the PoIs that need to be transmitted. The camera utilizes the HTTP protocol and creates request sessions to send POST requests to the server. The server returns the position information for PoIs of the next frame to the camera in JSON format. The ViT detector is implemented with Python and trained using MMTracking [35] framework on the edge server. The PPS is implemented with C++ on the camera. Figure 7 shows our experimental platform.
@@
Figure 7: Snapshot of our experimental evaluation hardware platform, including an edge server with GPUs and an NVIDIA Jetson AGX as the camera.
V Evaluation
@@
V-A Experimental Setup
V-A1 Performance Metrics
@@ V-A2 Datasets
$\bullet$ MOT17-Det [26]: It contains moving or stationary street views with pedestrian bounding box labels from indoor and outdoor settings across 11 distinct regions. The resolutions of these videos are either $640\times 480$ or $1920\times 1080$. We split the initial 75% of each video sequence to constitute our training set, with the remaining 25% serving as the validation set.
$\bullet$ AI City Challenge 2022 (AIC22) [27]: AIC22 is a real-world multi-camera traffic dataset including 42,683 frames and 88,820 vehicle bounding box labels. It contains video footage from a total of 59 surveillance cameras situated at five traffic intersections. We select videos from three intersections to constitute our training set while setting the videos from the remaining intersections as the validation set.
V-A3 Arena’s Configurations
We employ Faster R-CNN [32] as the two-stage detector framework and ViT-small as the backbone, which features a 12-layer transformer encoder and $D=384$. MFR is a one-layer transformer decoder that comprises merely 2.07M (5% of total) parameters and requires approximately 17.16 (1.9% of total) GFlops. The backbone pre-trained weights are from TinyMiM. Unless specifically stated, $p=0.9$, and $F=200$. For the expanded bbox margin $m$ and optical flow threshold $\beta$, we set them to $1$ and $10$ for the MOT17Det dataset, and $3$ and $1.5$ for the AIC22 dataset.
V-A4 Baselines
@@
$\bullet$ Faster R-CNN Detector [32] (FRD): Similar to FD, but the edge server employs a Faster R-CNN detector with ResNet-50.
@@
$\bullet$ DDS [5]: DDS transmits low-quality videos to the edge server to identify RoIs, which are fed back to the camera. The camera subsequently transmits these regions in high quality for second-round inference. We simulate different qualities by adjusting the resolution.
@@
$\bullet$ Reducto [28]: Reducto filters out frames with minimal impact on the accuracy by comparing low-level video feature differences between two adjacent frames at the camera to reduce bandwidth consumption. We employ Edge feature and utilize training sets to profile the optimal threshold.
@@
$\bullet$ ELF [10]: ELF uses attention-based LSTM to predict the bounding box shifts frame by frame. It then employs a frame partitioning method to offload the region proposals to multiple edge servers. We utilize the official implementation and configure it for three servers with parallel offloading and inference.
V-B Overall Performance of Arena
Figure 8: The normalized bandwidth usage of different methods on two datasets.
@@ V-B Overall Performance of Arena
Figure 9: The accuracy of different methods on two datasets. Arena can maintain accuracy losses within 1% and 4%.
@@
Figure 10: The average end-to-end latency per frame of different methods on two datasets. End-to-end latency includes a breakdown of preprocessing, transmission, and inference time.
@@
In Figure 8 and Figure 9, we report the normalized bandwidth usage and accuracy of different methods on two datasets. Figure 10 displays the end-to-end latency of each method, including a breakdown of preprocessing, transmission, and inference time. Among these baselines, Arena achieves the best trade-off among bandwidth, latency, and accuracy in both datasets. FD and FRD always transmit the full frame for inference, representing the upper bound for bandwidth consumption. They achieve the highest accuracy on their models but also result in the longest end-to-end latency. DDS achieves high bandwidth efficiency by transmitting only a few regions in high quality but inevitably incurs additional delay due to two rounds of inference. Reducto struggles to adapt to object detection tasks, as removing too many frames significantly reduces its accuracy. ELF maintains low bandwidth consumption and end-to-end latency, but this is contingent upon the premise of parallel offloading and inference by three servers. In contrast, Arena substantially saves bandwidth usage by pruning patches in non-keyframes, thus reducing both transmission delay and the model inference time. Meanwhile, Arena utilizes historical frame information to reconstruct features, minimizing the impact on accuracy. And Arena integrates AKIS into the end-to-end inference in parallel, which does not affect the end-to-end inference latency. Specifically, Arena achieves a maximum acceleration of end-to-end latency by 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × in two datasets, using only 47% and 31% of the bandwidth, while maintaining accuracy losses within 1% and 5% as compared to FD, respectively.
@@
Figure 11: Visualization of Arena on two videos. In these two scenes, with a frame interval of 5, $m$ is set to 1 and 3 for MOT17 and AIC22, respectively, $p=0.9$, and $F=200$. Only the red patches in non-keyframe are used for transmission and inference.
@@
Figure 12: Heatmaps of patches identified as PoIs, where darker areas indicate a higher frequency of offloading to the edge server.
@@
Moreover, Figure 13 presents the cumulative distribution function (CDF) charts of the proportion of PoI extraction in non-keyframes across various scenes. Taking the AIC22 dataset as an example, PoIs are extracted in a more granular and sparse manner, with about 80% of non-keyframes requiring the transmission of less than 25% area of the frame. This is consistent with our findings in the motivation (Sec. II-B). The results above reveal that Arena can effectively exploit the extensive redundancy present in the video to achieve end-to-end acceleration.
Figure 13: The cumulative distribution function (CDF) of PoI proportion of all non-keyframes.
Figure 14: Variation of AKIS-based keyframe interval in different scenes in the two datasets.
In Figure 14, we show the effect of AKIS algorithm on keyframe interval in a variety of scenes, as implemented in Arena. In the MOT17Det dataset, the upper limit of the keyframe interval is set to 15, and it is evident that AKIS is capable of adapting the keyframe interval in a manner that is conducive to inference in a range of scenes. For example in certain scenes exhibiting gradual alterations or a limited number of objects—such as “MOT17-02” or “MOT17-04”, AKIS will increase the inference interval in a manner that is adaptive to the circumstances. Therefore, AKIS serves to diminish the number of inference operations required for the transmission and inference of complete keyframes, while concurrently ensuring the maintenance of accuracy. Furthermore, this strategy enables the reduction of both the inference time and the bandwidth required for the process. In more complex scenes, such as “MOT17-09” and “MOT17-11”, AKIS will reduce the keyframe interval in order to ensure accuracy. A comparable phenomenon can be observed in the AIC22 dataset. The results demonstrate that the AKIS algorithm, as proposed in Arena, is capable of adaptively modifying the keyframe interval in accordance with scene alterations. This enables the attainment of a balance between accuracy and bandwidth in real-time scenes.
TABLE II: The results of whether AKIS is applied to Arena on the two datasets, and the comparison with the fixed keyframe interval.
Datasets MOT17Det AIC22
AKIS Keyframe Interval mAP@0.5 BandWidth Usage mAP@0.5 BandWidth Usage
Yes - 0.813 47% 0.854 31%
No 1 0.82 100% 0.902 100%
No 5 0.810 53.5% 0.853 34%
No 10 0.804 47.3% 0.829 25.4%
Figure 15: The impact of different sampling rates on accuracy.
V-C Ablation Study
In order to compose Arena, we propose several optimization adjustments for accuracy and bandwidth, 1) AKIS algorithm, 2) PPS mechanism. We conduct ablation studies on these modulers to demonstrate their effectiveness for Arena.
AKIS algorithm. We first compared the results on both datasets with and without the AKIS algorithm implemented in Arena. In Table II, compared to the keyframe interval fixed as 5, the AKIS algorithm reduces bandwidth usage dramaticly while maintaining a slight increase in accuracy. In consequence, it can be demonstrated that the AKIS algorithm can be effectively deployed in Arena, ensuring reconciliation of accuracy and bandwidth consumption while allowing for adaptive adjustment of the keyframe interval based on the complexity of the scenes.
@@ V-C Ablation Study
PPS mechanism. We illustrate the impact of sampling rates within high probability regions on accuracy in Figure 15. For MOT17Det, accuracy remains relatively stable when the probability $p$ is within the range of 0.8 to 1.0. Similarly, for AIC22, a $p$ value exceeding 0.9 consistently keeps high accuracy. This indicates Arena’s strong capability to compensate for the input. Concurrently, in comparison to the transmission of the entire PoI region ($p=1$), the PPS mechanism is capable of effectively reducing the consumption of transmission bandwidth within the overall video stream, while maintaining accuracy.
V-D Model Analysis
@@
In this subsection, we compare the model complexity of using an ViT detector with other commonly used detectors. As shown in Table III, we benchmark the GMACs and GFlops at a resolution of $800\times 1333$ using calflops [^1] https://github.com/MrYxJ/calculate-flops.pytorch and report the frame per second (FPS). ViT-small is employed as the backbone instead of the ViT-base version in Arena because it can achieve better real-time performance while already maintaining good accuracy. When processing the videos in the dataset, Arena operates with a computing load below 25% or 50% versions in the non-keyframe inference phase. This is because the proportion of PoI in most scenes is below 50% and even 25% as shown in Figure 13.
@@
Finally, we report the model precision in Table IV. Though Faster R-CNN offers the fastest inference speed, it requires the transmission of full frames, thus slowing down the end-to-end latency. Other models [36, 37] using the detection transformer [38] architecture, despite achieving commendable accuracy, fall short in real-time performance.
@@
TABLE III: Comparison of model complexity with 800 $\times$ 1333 resolution. Here we compare the inference speed (FPS) and complexity (GFlops) between widely used object detectors with different ratios of patch proportion.
@@
Faster R-CNN [32] ResNet50 41.123 214.648 430.087 47.47
Deformable-DETR [36] ResNet50 39.823 203.641 408.499 2.73
DINO [37] ResNet50 47.265 286.45 574.363 2.07
@@
TABLE IV: Model Precision
@@
Faster R-CNN [32] ResNet50 0.791 0.587 0.865 0.620
Deformable-DETR [36] ResNet50 0.805 0.630 0.880 0.648
DINO [37] ResNet50 0.819 0.632 0.876 0.656
VI Related Work
In this section, we briefly review the latest work in video analytics and also discuss efforts related to accelerating inference with Vision Transformers.
VI-A Real-Time Video Analytics Systems
@@
Real-time video analytics is a computationally intensive task. Substantial effort [5, 6, 7, 8, 9, 10, 11, 39, 40, 4] focuses on enhancing the real-time performance of systems in various ways. For example, to reduce the transmission data, server-driven methods [5, 41] allow cameras to send low-quality videos to the edge server. The edge server then identifies RoIs and provides the camera with their position feedback. In the second transmission, only these RoIs, encoded in high quality, are sent to the server to be processed. Another line of work [39] adopts a lightweight approach that directly extracts RoIs at the cameras and transmits only these regions to the edge server for inference. Lv et al. [7] propose a feedback mechanism for detecting RoIs, which involves extracting RoIs from a frame and stitching them together onto a single image for inference. This approach effectively reduces both bandwidth consumption and latency. Liu et al. [8] use optical flow to identify regions of interest in adjacent video frames and mask the remaining areas to reduce bandwidth consumption. Additionally, some cloud-edge collaborative systems [42, 43] enhance system efficiency by offloading entire or partial computational tasks to the cloud. Yang et al. [40] develop a strategy that selects whether to offload computing tasks to the Edge or the Cloud based on the complexity of the video content. Peng et al. [24] partition object areas from high-resolution videos into patches and offload them to a serverless platform for batch processing.
However, our work is the first to fully leverage characteristics of vision foundation models and achieve acceleration both on computing and transmission in the video analytics system.
VI-B Efficient Methods for Vision Transformer
@@
Utilizing the MSA mechanism in ViT leads to computational complexity that typically scales quadratically with the number of tokens. Consequently, recent efforts have focused on reducing the number of tokens to improve the efficiency of ViT. These methods primarily fall into two categories: token pruning [19, 20, 44, 45] and token merging [46, 22].
@@
Token pruning adjusts the amount of tokens based on specific methodologies. Just like what has been done to CNN-type architecture pruning in the past few years, token pruning is applied to accelerate ViT. Rao et al. [19] introduces a trainable prediction module (i.e., MLP) that leverages both local and global token information to create a binary mask, which determines the token count for the next layer. In order to retain some of the information present in the pruned tokens, which may contain contextual data regarding factors such as location and background, several methods [20, 44, 45] suggest combining the pruned tokens into a single token. Liang et al. [20] categorize tokens into attentive tokens and inattentive tokens based on attention mechanism scores. Inattentive tokens are fused together and jointly inputted to the next layer network along with attentive tokens. Long et al. [44] simultaneously consider both token diversity and importance by performing mering on inattentive and attentive tokens, thereby enhancing accuracy. Wei et al. [45] merge the pruned tokens with the most similar attentive tokens to retain their own information. Most merging-based methods rely heavily on clustering algorithms, such as K-Means [46], K-Medoids, and density-peak clustering with K-Nearest neighbors to cluster and merge tokens. However, these methods present significant challenges in terms of convergence. Bolya et al. [21] propose a bipartite soft matching algorithm to cluster tokens and subsequently merge similar tokens, thus making it easier to converge.
@@
These ViT acceleration techniques by token reduction are not suited for Arena. First, they cannot reconstruct complete feature maps, which are necessary for dense prediction tasks. Second, they require the entire frame as input. Our proposed Arena, thinking out of the box, effectively adapts efficient ViT for video analytics with its advantages.
VII Conclusion